# Multilingual Performance of a Multimodal Artificial Intelligence System on Multisubject Physics Concept Inventories

### v1

@@ II Data set
EN, TH, ZH | THERM |
| TOAST | Test of Astronomy Standards | Astronomy content knowledge (gravity, electromagnetic radiation, fusion and formation of heavy elements, evolution of the universe, star and stellar evolution, evolution and structure of the solar system, seasons, scale, yearly patterns, daily patterns, moon phases) | [121] | vf | 44 [121] | EN, JA | AST |
| TUG-K | Test of Understanding Graphs in Kinematics | Mechanics content knowledge (kinematics, graphing) | [27] | 2.6 | 59 [122] | AR, DE, FI, FR, HE | MECH |
| 3.0-4.0 | EL, EN, ES, PT, SV, UK, ZH | ||||||
| TUV | Test of Understanding of Vectors | Mathematics content knowledge (magnitude, direction, components, unit vector, addition, subtraction, multiplication, dot and cross product) | [123] | 1 | 68 [123] | AR, EN, ES | MATH |
@@ III.1 Data preparation
The questions from the concept tests were captured using screenshots like the ones shown in Fig. 1. If a question had multiple parts referring to the same scenario or each other, these were combined in one image; at times, this required manual photo-editing to close page breaks. Questions that assessed confidence levels (“How sure are you of this answer?”) or those with free-response answers were skipped. The tests in Tables 3–6 resulted in 3,662 images.
The solution keys were transcribed from PhysPort, where we transformed the various option values such as B), 2., b., ), and non-Arabic, non-Latin, non-Greek characters into lower-case Latin characters. For tests that gave partial credit for certain incorrect answers if previous answers were incorrect (“follow-up mistakes”), we simplified the evaluation by only giving credit for correct answers. Also for tests that assigned different point values for different questions, we only considered correctness. In addition, for each question, we coded if it was text-only, or if it included an image, graph, or scenario illustration.
@@ III.2 AI processing
We have been using GPT-4o [25] Version 2024-08-06 via Microsoft Azure AI Services [124] at ETH Zurich. The university’s contract includes provisions that any data submitted will not be used for training purposes; this provision is crucial to avoid compromising the confidentiality and validity of the concept tests.
The screenshots were submitted via the deployment’s API in base64-encoding as shown in Fig. 2. As role, we used the text shown in Fig. 3, and as prompt the text shown in Fig. 4. To facilitate further processing and evaluation of the answers, we demanded the structured JSON output schema shown in Fig. 5. We included the “explanation” field in the output structure to capture the AI system’s reasoning steps, also referred to as Chain-of-Thought (CoT) [10]. This allowed us to glean some key insights into the system’s language behavior.
Each screenshot was independently evaluated three times at a temperature of 0.7 (which is the default in the chat clients), and the resulting three problem structures were combined into a solution array. This resulted in output like the one shown in Fig. 6. Altogether, we obtained 14022 solutions for 4,674 questions (1,498 for English and 3,176 for non-English language inventories).
As Fig. 6 illustrates, the role and prompt we used can result in mixed-language outputs. While this issue could probably have been avoided by translating the role and prompts into the language of the tests, doing so for languages using non-ASCII characters would likely have made the mapping between the prompt and the fields in the JSON structure unreliable. Colloquially speaking, “the AI was allowed to think in the language of its choice.”
@@ III.3 Analysis methods
The answer choices found in the JSON structures were normed to the same lower-case Latin characters as the solution keys. In cases where the AI did not provide a valid answer (e.g., where it claimed that there was no correct answer specified), the answer was counted as incorrect. Some concept inventories had images or scenarios labeled with Roman characters and answer choices such as “A) I, B) II, C) III, D) IV, E) none of the above;” in this case, if the AI picked “IV,” this was manually converted to “d.” Another source of possible error were numbered scenarios within numbered multipart problems; in this case, the AI at times ignored the problem numbers provided in the prompt and instead used the scenario numbers; this was fairly obvious during the evaluation, but had to be fixed manually.
We counted each answer in an inventory — three answers per question — as either correct or incorrect, and considered the percentage of correct answers as performance measures. For each answer, we coded the language of both the test description and the answer explanation, primarily using langdetect [125] and manual determination in some cases. Responses were categorized as fully in the language of the test, fully in English, or mixed. The most common mixed scenario occurred when the problem description was still in the language of the test, but the explanation was in English. However, there were also rare cases where the language switched mid-stream for either of those.
@@ III.4 Use of AI
While obviously being the subject of this study, AI (GPT-o1 [126]) has also been used for the following aspects of the study:
Initial drafts of analysis programs in R and Python.
Exploratory data analysis.
LaTeX formatting of manuscript components.
Improving the grammar and readability of manuscript passages.
@@ IV Results
IV.1 Scores by inventory and language
@@
Table 7 shows the scores on each of the inventories in each available nominal language (as “nominal,” we denote the language in which the inventory was presented). One immediate observation is the uneven coverage of assessments across languages. Many inventories are only available in a handful of languages, making broad comparisons challenging. However, when data is present, certain patterns do emerge. Some assessments, such as FCI, have results for numerous languages, allowing a more comprehensive look at language-to-language differences.
In the FCI, performance ranges from as low as 20% in Punjabi and 22% in Tamil to as high as 74% in Portuguese and Polish, suggesting that the AI’s proficiency can vary significantly depending on the language used. Since the FCI uses five answer options, a 20% performance is equivalent to randomly picking an answer. An in-depth analysis of the FCI answers for Punjabi revealed that GPT-4o started hallucinating, generating output about uniform magnetic fields and charged particles. In Tamil, the explanations and conclusions were sometimes correct, but then GPT-4o picked the wrong answer letter. Other assessments, like TUV, ADT, and FMCE, show relatively consistent scores in the 60–80% range wherever data is available.

### v3

@@ II Data set
Chinese, English, Thai | THERM |
| TOAST | Test of Astronomy Standards | Astronomy content knowledge (gravity, electromagnetic radiation, fusion and formation of heavy elements, evolution of the universe, star and stellar evolution, evolution and structure of the solar system, seasons, scale, yearly patterns, daily patterns, moon phases) | [150] | vf | 44 [150] | English, Japanese | AST |
| TUG-K | Test of Understanding Graphs in Kinematics | Mechanics content knowledge (kinematics, graphing) | [34] | 2.6 | 59 [151] | Arabic, Finnish, French, German, Hebrew | MECH |
| 3.0-4.0 | Chinese, English, Greek, Portuguese, Spanish, Swedish, Ukrainian | ||||||
| TUV | Test of Understanding of Vectors | Mathematics content knowledge (magnitude, direction, components, unit vector, addition, subtraction, multiplication, dot and cross product) | [152] | 1 | 68 [152] | Arabic, English, Spanish | MATH |
We cannot exclude the possibility that some of these inventories appeared in the text corpus used to train the model, which could be seen as “teaching to the test.” However, GPT’s training corpus is closed-source and its contents remain mostly speculative with only very few hints given [153]. While popular inventories like the FCI might have been included, it is less likely that many of the more obscure inventories were present. Furthermore, in the scientific literature, the solution keys are typically provided separately from the inventory items. This means that the model is unlikely to associate correct answers to their corresponding questions on the basis of proximity in the training data.
@@ III.1 Data preparation
The items from the concept inventories were captured using screenshots like the ones shown in Fig. 1. If an item had multiple parts referring to the same scenario or each other, these were combined in one image; at times, this required manual image-editing to close page breaks. The inventories in Tables 2–5 resulted in 3,662 separate image files that were submitted to the model.
The solution keys were transcribed from PhysPort, where we transformed the various option values such as B), 2., b., ), and non-Arabic, non-Latin, non-Greek characters into lower-case Latin characters for easier automatic processing.
For the MUQ, which assigns partial credit for incorrect answers, we simplified the evaluation by only giving credit for correct answers; this would under-estimate the model’s performance. We did the same for BEMA, which has conditional grading rules for four out of 31 items, depending on answers from earlier items. Here, for two items, this simplification is in the model’s favor, while for the other two items, it is to the model’s disadvantage. In the FTGOT, we excluded the items assessing certainty in the answer and compared results for a two-tier test version. Finally, we skipped the free-response items that were included in some of the inventories, for example to explain reasoning, as those were not scored in the original inventories.
For each individual item, we coded if it was text-only, or if it included an image, graph, or scenario illustration. Additionally, we coded if consulting the image was necessary for correctly answering the question (required image), or if all the required information was already present in the text (unneeded image). These codes will be used in analyses of subclasses of problems in Section IV.D.
@@ III.2 AI processing
We used GPT-4o [32] Version 2024-08-06 via Microsoft Azure AI Services [154] at ETH Zurich. The university’s contract includes provisions that any data submitted will not be used for training purposes; this provision is crucial to avoid compromising the confidentiality and validity of the concept inventories.
The model was prompted to extract from each submitted image the number and the text of the inventory item, followed by written-out reasoning steps (explanation), and finally the letter option corresponding to its selected answer. In cases where multiple items appeared in the same screenshot, the model was instructed to repeat this process for each item. To process such outputs effectively, we had to prompt the model to return structured outputs in the form of a JSON schema. We found that prompting for structured outputs in languages other than English was unreliable and hindered the type of analysis we aimed to perform. As a result, we decided to keep the prompt in English. Each submitted image, however, contained text in one of the different languages of the concept inventories. The API call and prompts used in this study are available in Appendix A.
LLMs are probabilistic systems, and thus responses to the same prompts vary. To obtain some statistics, each screenshot was independently submitted three times, and the resulting three outputs were combined into a solution array.
Altogether, we obtained 14,022 solutions for 4,674 items (1,498 for English and 3,176 for non-English language inventories).
@@ III.3 Analysis methods
The LLM’s answer choices were normed to the same lower-case Latin characters as the solution keys. In cases where the AI did not provide a valid answer (e.g., where it claimed that there was no correct answer specified), the answer was counted as incorrect. Some concept inventories had images or scenarios labeled with Roman characters and answer choices such as “A) I, B) II, C) III, D) IV, E) none of the above;” in this case, if the AI picked “IV,” this was manually converted to “d;” this happened for 0.4% of the responses. Another source of possible error were numbered scenarios within numbered multipart problems; in this case, the AI at times ignored the problem numbers provided in the prompt and instead used the scenario numbers; this was fairly obvious during the evaluation, but had to be fixed manually. In 0.6% of the cases, the LLM provided no valid response; these were counted as incorrect.
We counted each answer in an inventory — three answers per item — as either correct or incorrect, and considered the percentage of correct answers across each inventory and language as the performance measure for that test and language. For each answer, we coded the language of both the inventory description and the answer explanation, primarily using langdetect [155] and manual determination in some cases. Responses were categorized as fully in the language of the test, fully in English, or mixed. The most common mixed scenario occurred when the problem description was still in the language of the test, but the explanation was in English. However, there were also rare cases where the language switched mid-stream for either of those. We refer to this behavior as language switching.
@@ III.4 Use of AI
While obviously being the subject of this study, AI (GPT-o1 [41]) has also been used for the following aspects of the study: initial drafts of analysis programs in R and Python, exploratory data analysis, LaTeX formatting of manuscript components, and improving the grammar and readability of manuscript passages.
@@ IV Results
IV.1 How does GPT-4o perform across different physics concept inventories in English?
Notably, English was the language with the most concept inventories: 53 out of 54 tested concept inventories were available in English. It is also the only language that has inventories across all the subject categories. In Table 6, the English results for individual inventories stand out as the most populated column. The average performance across all inventories in English is 71.1%, with results on individual inventories ranging from as low as 27% (FTGOT) to as high as 97% (DS and STPFaSLlo). Figure 2 shows the distribution of GPT-4o’s performance across the different concept inventories in English.
Examining the performance within specific subject categories, it is noticeable that for some subjects, the performance varies widely across inventories. As Fig. 3 shows, this includes outliers in the 30% range (that is, only slightly better than randomly picking answer options):
Within quantum physics (QP), this outlier is the Quantum Mechanics Visualization Inventory (QMVI) at 32%. This inventory heavily focuses in graphical visualizations of wave functions.
Within optics (OPT), the outlier is the Four-tier Geometrical Optics Test (FTGOT) at 26%. This inventory deals with ray optics, another graphical visualization topic.
However, the extreme values on individual inventories mostly average out — except in categories with few inventories (e.g., Relativity and Laboratory skills).
The AI performs best in Thermodynamics (85.2%), followed by Astronomy (80.4%), Reasoning (78.7%), and Relativity (75.0%). Inventories in Astronomy tend to focus on factual knowledge, while those in Thermodynamics, Reasoning, and Relativity tend to use exact language. The AI’s performance is weaker in Laboratory skills (35.0%), which include strategies for data collection and analysis.
IV.2 How does language influence the performance of the AI system?
@@
Table 6 also shows the scores on each inventory in its respective nominal language (we use “nominal” to denote the language in which the inventory was presented). One immediate observation is the uneven coverage of assessments across languages. Many inventories are available in only a handful of languages, making broad comparisons challenging. Some assessments, however — most notably FCI — are available in numerous languages.
In the FCI, performance ranges from as low as 20% in Punjabi and 22% in Tamil to as high as 74% in Portuguese and Polish, suggesting different performance across languages. Since the FCI uses five answer options, a 20% performance is equivalent to random guessing.